INTERACTIONS IN SYMMETRIC TOP MOLECULES BETWEEN VIBRATIONAL POLYADS: ROTATIONAL AND ROVIBRATIONAL SPECTROSCOPY OF LOW-LYING STATES OF PROPYNE

\underline{HOLGER S. P. MÜLLER}, I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany; PETR PRACNA, Academy of Sciences of the Czech Republic, 18223 Praha 8, Czech Republic; VELI-MATTI HORNEMAN, 90014 University of Oulu, Finland; STEFAN KLEE, Justus-Liebig-Universität Gießen, 35392 Gießen, Germany.

A large body of very accurate (mostly 10–20 kHz) rotational transitions within the 10 μm tetrad of the strongly prolate symmetric rotor propyne, H_{3}C–C≡CH, had been recorded in order to analyze some weaker resonances.^{a} The analysis, in particular for higher K transitions, indicated new problems: Interactions of members of the 10 μm tetrad (v_{5} = 1, v_{9} = v_{10} = 1, v_{10} = 3, and v_{8} = 1) with those of the 15 μm dyad (v_{9} = 1 and v_{10} = 2). More than 300 rotational transitions within v_{10} = 1 (E_{vib} = 330.9 cm^{−1}) have been recorded with J′′ ≤ 53 and −11 ≤ K · l_{10} ≤ +16 for a more systematic investigation into the lower excited vibrational states of propyne. Even the inclusion of many high order terms prevented the highest K transitions to be fit within experimental uncertainties, in particular those having K · l_{10} ≤ 0. Inclusion of ν_{10} IR transitions from a reanalysis of the spectrum that was used in the study of the 10 μm tetrad^{a} confirmed these findings. It turned out that the first order energies of the K · l_{10} = –12 levels coincide to within 1 cm^{−1} with those of the overtone v_{10} = 2, K = 12 of the l_{10} = +2 substate. A weaker Coriolis interaction occurs between v_{10} = 1, K · l_{10} = –11 and v_{9} = 1, K · l_{9} = +10.^{b} Meanwhile, the investigation of v_{10} = 1 has been completed, and those of v_{9} = 1 and v_{10} = 2 have begun. Fermi-type interactions occur between the states v_{10} = 2 and 3 and between v_{9} = 1 and v_{9} = v_{10} = 1. Among the Coriolis resonances, the one between v_{10} = 2, l_{10} = –2 and v_{5} = 1 permitted intervibrational transitions to be detected in the submillimeter region. These interactions are expected to improve A, D_{K}, etc. for v = 0. Selected details of our ongoing analyses will be presented. In addition, the relevance of these types of interactions for other symmetric top molecules will be discussed.

aP. Pracna, G. Graner, J. Cosléou, J. Demaison, G. Wlodarczak, V.-M. Horneman, and M. Koivusaari, J. Mol. Spectrosc. 206, (2001) 150–157; H. S. P. Müller, P. Pracna, et al., unpublished

bH. S. P. Müller, P. Pracna, and V.-M. Horneman, J. Mol. Spectrosc. 216, (2002) 397–407